BOUNDARY DETERMINATION OF A RIEMANNIAN METRIC BY THE LOCALIZED BOUNDARY DISTANCE FUNCTION

GUNTHER UHLMANN AND JENN-NAN WANG

Abstract. In this paper we give a reconstruction procedure for the Taylor series of a Riemannian metric on the boundary in boundary normal coordinates from the localized boundary distance function.